OpenStudy (anonymous):

Given y=arcsin(-3/7), find cot y. An explanation of how to do it would be very helpful. Thanks

OpenStudy (anonymous):



OpenStudy (anonymous):

you want the cotangent of that angle, so all that is missing is the "adjacent' side which you find via pythagoras

OpenStudy (anonymous):

\(\sqrt{7^2-3^2}=\sqrt{49-9}=\sqrt{40}=2\sqrt{10}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

now cotangent is "adjacent over opposite" so you are just about done except for the sign of the answer

OpenStudy (anonymous):

Got it, it makes more sense now. so the answer is (-2√10)/3 right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw